John spent $352 on fencing for his dog pen. The fencing cost $11.00 per foot. He built a rectangular pen 10 feet in length. If he had planned better, he could have increased the area avalible to his dog by a maximum of how many square feet?
Answer by mananth(16949)

$352 for the fence @ $11 . foot.
so total fencing = 352/11= 32 feet.
..
Perimeter P = 2(L+W)
32= 2(L+W)
32/2 = L+W
16= 10+W
W= 16-10
W= 6 feet.
..
The area of the pen = 10*6 = 60 sq. ft.
we know square gives maximum area.
so 8*8 =64
He could have gained 4 sq.ft